OpenStudy (anonymous):

Solve y over negative 2 + 5 = 13. a.-16 b.16 c.-36 d.36

OpenStudy (anonymous):

@e.mccormick

OpenStudy (akashdeepdeb):

y/2 + 5 = 13 So y/2 = 13-5 [Subtract 5 from both sides] y/2 = 8 y = 8*2 [Multiply 2 to both sides] y = 16 Understood? :)